OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the points (-2, 6) and (-1, 4).

OpenStudy (anonymous):

Use slope formula m = y2-y1/x2-x1 then use point-slope form of line: y-y1 = m(x-x1)

OpenStudy (anonymous):

I got y=-2x+0 but I know that's not right